OpenStudy (anonymous):

How would you differentiate f(x)= [ e^arctanx] \ 1+x^2? I used the Quotient rule as u= e^arctanx and v=1+x^2 so, I have f ' (x)= [ (1+x^2) d\du e ^arctanx - e^arctanx * d\dv(1+x^2) ] \ (1+ x^2)^2 Then I used the composite rule , with g(u)= e^u g'(u)= e^u u= f(x) = arctan x f'(x)= 1\ 1+x^2 so f'(x) = [ (1+x^2)*(e^arctanx)*(1\1+x^2)- e^arctanx*(2x) ] \ (1+x^2)^2 f'(x)=[ e^arctanx - e^arctanx * (2x) ] \ (1+x^2 )^2 f'(x)= -2x \ (1+x^2 ) Is this ok? Please help, many thanks !

OpenStudy (amistre64):

im thinking logs ... f(x)= [ e^arctanx] \ 1+x^2 log(f) = log ([ e^arctanx] \ 1+x^2) log(f) = log (e^arctanx) - log(1+x^2) log(f) = arctanx - log(1+x^2) f'/f = (arctanx)' - 2x/(1+x^2) f' = f [(arctanx)' - 2x/(1+x^2)]

OpenStudy (amistre64):

arctan derivative is what ... 1/1+x^2 if memory serves?

OpenStudy (amistre64):

im thinking then .... e^(arctanx) * (1-2x) f' = ----------------- (1+x^2) ^2

OpenStudy (amistre64):

i think you lost track of the quotient rule along the way, which is easy to do since it tends to be the most convoluted rule

OpenStudy (amistre64):

with g(u)= e^u g'(u)= e^u ^^^^^^^ chain rule u' e^u

OpenStudy (anonymous):

Thank you !!! :)